Subject: APO 3.5m users committee minutes, May 2014

From: Suzanne Hawley

Submitted: Wed, 28 May 2014 22:43:34 -0700 (PDT)

Message number: 1332 (previous: 1331, next: 1333 up: Index)

 		APO 3.5-m Users Committee Phonecon, 5/19/14

Attending:  Suzanne Hawley, Mark Klaene, Rachel Kuzio de Naray, Bill 
Ketzeback, Bruce Gillespie, Rene Walterbos, Russet McMillan, John Bally, 
David Bowen (for Michael Strauss), John Wisniewski, Joanne Hughes Clark, 
Al Harper


User feedback and comments from institutional representatives:

  o Oklahoma - John Wisniewski said that a recent student trip went well.

  o Colorado - John Bally had nothing new to report.

  o Georgia State - Rachel Kuzio de Naray said that the GSU attendees very much 
enjoyed the 20-yr anniversary celebration.

  o Princeton - David Bowen said Princeton is happy.

  o Seattle Univ. - Joanne Hughes Clark reported that her recent time trade 
with WIYN went well; she used the pODI and felt that is was only a bit 
more complicated to use than SPIcam.  Her next run at WIYN will be remote.

  o Chicago - Al Harper had nothing new to report.  He enjoyed the trip to
APO and the 20th anniversary celebration.

  o NMSU - Rene Walterbos had nothing new to report.

  o Washington - Suzanne Hawley noted that there is a monitoring program 
using evening twilight for several weeks this quarter, and one user has 
expressed concern about obtaining their calibration data.  The schedule 
was written to minimize the effect on other users as much as possible, but 
we are interested in hearing suggestions from users about additional ways 
to mitigate.


Discussion of telescope/instruments report:

Mark highlighted the recent 20-year anniversary event held at APO, which 
was well-attended and enjoyed by all.  Also, the installation and 
troubleshooting of the new drive amplifiers was a major activity, and 
should fix long-standing problems we have had over the years with 
operation at the higher ambient air temperatures in May and June.  The 
rest of his comments were from statements contained in the written report, 
which follows:


 	      3.5-m Telescope and Instruments Highlights
 	                04/10/14 through 05/15/14

0) Overview

Generally another productive period with variable cloudiness, little 
precipitation and with typical spring wind and dust events. Current forest 
fire danger is now "Very High".  We completed the 30/20 year ARC/3.5m 
anniversary events on site with approximately 80 in attendance. There was 
a 100-acre forest fire 3 miles west of the observatory in April.  While it 
was a bit disconcerting, the observatory was not threatened and 
observations continued with minor interruptions.

1) Telescope

Telescope has been working well except for pointing issues.  These issues 
came about during the altitude amplifier upgrade work.  Some observations, 
especially with Echelle were affected a lot, but most were able to be 
salvaged with only small amounts of time lost.  It took a couple attempts 
over several days and nights at fixes and pointing corrections to isolate 
and correct the problem.  While we expected some issues with the amplifier 
swap the problem we saw was not expected and still somewhat unexplained.

Note that the new amplifiers should avoid lost time seen in past years due 
to overheating in late May and June when ambient temperatures are high.

The new TCC development is also progressing nicely with continued testing 
on the telescope.

2) Instruments

DIS has been operating normally.

Agile is operational with no problems.

TripleSpec had no operational problems this period.

SPIcam variable dark current was addressed this period. The Gamma Vacuum 
ion pump and controller continues to show current spikes about once per 
day however baffling was installed to eliminate the associated ion glow 
from reaching the detector. This fix appears to have been successful based 
on tests to date.  We are continuing to monitor this for a while longer.

Echelle has been operational. We had one ion-pump high-current step that 
lasted about 12 hours.  It was determined that it was caused by a whisker 
on the pump and once that was cleared the current level dropped back to 
normal range.  No effect was seen on the detector.

NIC-FPS continues to operate in shared-risk observing mode due to 
occasional image corruption issue.  NFS less than 9 is operational.

GIFS is operational.  The small LN2 hold time designed in by GFSC 
continues to be a problem.  The TUI interface is ready and undergoing 
final user testing and documentation upgrades for the upcoming training 


The instrument development is proceeding on schedule.  The CCD has been 
ordered.  Final testing of the prototype shutter is ongoing.  One of the 2 
cylinders failed at a high stress concentration point and design changes 
have occurred. The optics for the final f/8 design has been reviewed and 
approved by the group.  Once the final report is published and reviewed a 
go-ahead is expected.

4) Summer shutdown

There will be 2 shutdowns scheduled this summer.  The first is from July 
7th to July 24th for M2 and M3 mirror re-coating, mechanical maintenance, 
electronic system upgrades and the majority of instrument servicing.  A 
second shutdown will be from August 8th to August 17th for the new TCC 


o GIFS training run - Ketzeback:

Bill said that the hands-on GIFS training will be held May 27-28.  This is 
the only training session that is planned for the instrument, and there 
followed an extended discussion on the topic of long-term user access to 
the instrument.  Anyone who missed this training opportunity and wants to 
use GIFS should contact Carol Grady.


o ARCSAT update - Hawley, Ketzeback:

Bill said that ARCSAT is now in shared-risk use mode and fully scheduled 
through September.  Both survey-cam and flare-cam are being used.  We have 
been learning the quirks and features of the new telescope control 
software, and appreciate the feedback from the users.  Suzanne added that 
we've just purchased a Johnson-Cousins filter set for ARCSAT.  Bill has 
taken over the main engineering responsibilities for ARCSAT while Joe 
Huehnerhoff is now focussing on ARCTIC.  Those interested in using ARCTIC 
in the fall should familiarize themselves with the documentation under the 
ARCSAT link on the APO homepage.


WIYN trades update - Hawley:

Suzanne said that there will be two ARC programs scheduled on WIYN during 
their 2014B semester, and two programs from WIYN PIs will be scheduled on 
the 3.5m.  We are also considering talking with other telescopes for 
possible reciprocal use arrangements.


o 20th anniversary recap - Klaene, Hawley, Walterbos

Rene said the celebration went very well, and the logistics all worked 
smoothly.He gave special thanks to Gretchen Van Doren for all her hard 
work, and to the APO staff for their many contributions to the successful 
event.  The weather also cooperated as the evening star party allowed the 
visitors a great view of the sy through the 3.5m eyepiece.  It was very 
rewarding to see many old friends return to the site, and to hear about 
the early days of the telescope design and construction, as well as 
numerous science presentations.  Good discussions were also held with 
potential new ARC partners who were present.  Bruce mentioned that the 
historical talks underscored the need to capture the history of the 
telescope before it becomes lost.



[open from previous meeting]:


[new from this meeting]:



Next meeting:  The next Users Committee phonecon will be on Monday 23 June 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent to 
the committee members during the preceding week.  Note that we will then 
skip July as is our custom during summer shutdown.  The following meeting 
will be held on Monday 18 August.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1332 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to